1.     An object is projected is projected at an angle 0o to the ground at a speed V m/s Neglecting the air resistance, show that the range of the projectile is

R=V^2 sin20/g



1
Expert's answer
2021-05-26T10:33:32-0400

"R=v_x\cdot t=v\cdot\cos20°\cdot t"


"v\cdot\sin20°-gt=0\to t=\frac{v\cdot\sin20°}{g}"


"R=v\cdot\cos20°\cdot \frac{v\cdot\sin20°}{g}\cdot2=\frac{v^2\sin40°}{g}" . Answer